x + y = -4
4x + y = -13
y + 3z = -10

Answer by stanbon(75887)

1) x + y = -4
2)4x + y = -13
3)y + 3z = -10
1) y=-4-x
2) 4x+(-4-x)=-13
3x=-9
x=-3
Then from 1) y=-4-x
y=-4+3=-1
3) -1+3z=-10
3z=-9
z=-3
So, x=-3, y=-1, z=-3
Cheers,
Stan H.